# Physics-informed Hamiltonian learning for large-scale optoelectronic property prediction

### v1

@@ Introduction
Relevant in this context, learning the Hamiltonian is challenging because it depends not only on a system’s atomic structure but also on the chosen representation, so there is no single unique mapping from structure to matrix elements. However, equivariant ML models respecting fundamental physical transformations, such as rotations and translations, ensure that predictions transform consistently under symmetry operations, and thereby significantly improve data efficiency and generalization. Specifically, equivariant neural networks (ENNs) were recently developed as particularly promising frameworks for Hamiltonian learning [17, 18, 19]. However, ENNs typically require large datasets for training, which can be a limiting factor from a computational perspective. Related to this limitation, existing ENNs for Hamiltonian learning usually rely on access to the real-space Hamiltonian represented in a consistent basis, which restricts their applicability to specific DFT frameworks requiring careful benchmarking [20].
@@ Results
Fig. 1 : Overview of the physics-informed Hamiltonian learning model and workflow with results for GaAs. a Illustration of pairwise and environment-dependent electronic interactions arising from structural fluctuations in a chemical system and their effect on the electronic structure captured via an effective Hamiltonian, $\hat{H}_{\mathrm{eff}}$, in the Hamster approach. It combines a physical model and environment-dependent machine learning terms. b Schematic visualization of the environment descriptor. c Validation error, calculated as mean absolute error (MAE) across all eigenvalues, for a Hamster model trained on an increasing number of training structures for GaAs at $400\text{\,}\mathrm{K}$. d Comparison of residuals of a pristine tight binding (orange) and the Hamster model (blue) with respect to density functional theory (DFT) data, averaged over k-points and 100 snapshots, at $400\text{\,}\mathrm{K}$. Dashed vertical lines indicate the valence band maximum (VBM).
@@
$\delta H^{\textbf{R}}_{ij}=\sum_{n=1}^{N_{\mathrm{p}}}h_{n}\exp\Bigl(-\frac{|\textbf{x}_{n}-\textbf{x}^{\textbf{R}}_{ij}|^{2}}{2\sigma^{2}}\Bigr{missing}),$ (4)
@@
where the sum runs over all atoms and orbitals, indexed by $k$, within a cutoff radius, $r_{\mathrm{cut}}$. The smooth cutoff function, $f_{\mathrm{cut}}(r)=\frac{1}{2}[\cos(\frac{\pi r}{r_{\mathrm{cut}}})+1]$, ensures a gradual decay of contributions from distant atoms.
To provide the model with information about the interaction type, we augment the descriptor vector with several structural features: the equilibrium distance between atoms $i$ and $j$, the angular orientation of their respective orbitals relative to the connecting vector, $\textbf{r}_{ij}$, and the atomic numbers $Z_{i}$ and $Z_{j}$. These quantities can either be kept fixed during atomic motion or updated dynamically, thereby providing the model with the flexibility to accommodate additional physical effects such as thermal lattice expansion. This design ensures that the ML model focuses on learning changes in the local environment, thereby improving data efficiency by reducing the complexity of the learning task. Finally, we note that the ordering of values within the descriptor vector is crucial, as inconsistent arrangements can violate symmetry properties of the Hamiltonian. To preserve these symmetries, we carefully construct the descriptor such that its structure is invariant under operations that should leave the Hamiltonian unchanged, in line with the principles employed by other equivariant models [15, 16, 31, 18, 13, 14].
@@
Fig. 2 : Training and descriptor analysis for CsPbBr3. a Dependence of training (dark orange) and validation (black) loss on the number of training structures. b Principal component analysis of the kernel support points for 12 training structures. In both panels, we compare two different cutoff radii ($r_{\mathrm{cut}}$) for the interaction range, namely 5.5 $5\text{\,}\text{\AA }$ and $6.2\text{\,}\text{\AA }$.
Fig. 3 : Transferability of the model across temperatures and large-scale calculations. a Band gap of CsPbBr3 computed with DFT (black) and Hamster (blue) for a $2\times 2\times 2$ supercell of CsPbBr3. Results obtained using Hamster for a $16\times 16\times 16$ supercell are shown as well (green). Thin vertical lines indicate the standard deviations. b Comparison of residuals of a pristine tight-binding (orange) and the Hamster model (blue) with respect to DFT data, averaged over $k$-points and 100 snapshots, for CsPbBr3 at $425\text{\,}\mathrm{K}$.
@@
The learning curves (see Fig. Physics-informed Hamiltonian learning for halide perovskitesa) show that the validation error plateaus after using around 10 training structures, reaching an accuracy for the eigenvalues that drops below $50\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$. While this level of accuracy is comparable to that reported in a recent study [21] using a neural-network scheme for elemental and binary semiconductor compounds, we emphasize that our approach needs far fewer training structures for the considerably more complex halide perovskite CsPbBr3, well below the few hundred typically required by neural-network-based Hamiltonian learning models [17, 31, 21]. Remarkably, for this material Hamster still generalizes reasonably well with as few as two training structures, remaining within $0.02\text{\,}\mathrm{e}\mathrm{V}$ of the training error, highlighting the markedly superior data efficiency of our approach. The mild overfitting at very low training set sizes, notable through the small gap between training and validation error, disappears as the training set increases.
@@
We can gain insight into the model’s increasing expressiveness when $r_{\mathrm{cut}}$ is increased by analyzing the structure of the descriptor space. To this end, we perform a principal component analysis (PCA, see Methods section) on the sampled descriptors for both cutoff radii, using twelve training structures (see Fig. Physics-informed Hamiltonian learning for halide perovskitesb).
@@
Before turning to large-scale simulations, we apply the model to MD trajectories of a $2\times 2\times 2$ supercell of CsPbBr3 for benchmarking against DFT data, using $r_{\mathrm{cut}}=$ 5.5 $5\text{\,}\text{\AA }$ in our model. Fig. Physics-informed Hamiltonian learning for halide perovskitesa compares the temperature-dependence of the band gap in CsPbBr3 as obtained with Hamster and DFT. Although the model was trained solely on structures from an MD trajectory at $425\text{\,}\mathrm{K}$, its predictions remain in close agreement with DFT also at higher temperatures. Across all three considered temperatures, the difference between the DFT-computed band gap and the value predicted by Hamster remains below $50\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$, and the MAEs of energy eigenvalues are $0.055\text{\,}\mathrm{e}\mathrm{V}$ ($425\text{\,}\mathrm{K}$), $0.056\text{\,}\mathrm{e}\mathrm{V}$ ($525\text{\,}\mathrm{K}$), and $0.058\text{\,}\mathrm{e}\mathrm{V}$ ($625\text{\,}\mathrm{K}$). Furthermore, we find that the fluctuations of instantaneous band gaps, gauged by the standard deviation at each temperature, are similar in DFT and Hamster. These results underscore the transferability of our model over a considerable temperature range within the cubic phase of CsPbBr3.
@@
A key advantage of our approach is that it enables simulations of system sizes far beyond the practical reach of explicit first-principles methods such as DFT. Modeling of larger systems can reveal distinct dynamical behavior and, by reducing finite-size effects, provide a more realistic description of the material. To explore this regime, we train an MLFF for CsPbBr3 (see Methods) and perform MD simulations for a $16\times 16\times 16$ supercell of CsPbBr3 consisting of 20480 atoms. Using the Hamster model, we subsequently predict Hamiltonians and compute band gaps by extracting eigenvalues at the VBM and CBM via the Lanczos algorithm as implemented in Arpack [43]. Figure Physics-informed Hamiltonian learning for halide perovskitesa shows that the mean value of the band gap at $425\text{\,}\mathrm{K}$ closely matches the one obtained for the smaller supercell. At the same time, the sizable band-gap fluctuations seen for the smaller supercells are almost entirely averaged out, yielding a much narrower distribution and a more realistic description of the material at finite temperature.
@@
We optimize the model parameters following the same procedure as for CsPbBr3 (see Methods section), again using $r_{\mathrm{cut}}=$ 5.5 $5\text{\,}\text{\AA }.$ However, for this material we do not include orbitals stemming from the cation and perform classical force-field MD calculations (see Methods section). When applied to an independent test set consisting of 100 MD snapshots of a $2\times 2\times 2$ supercell at $300\text{\,}\mathrm{K}$, the model reaches an MAE of $0.057\text{\,}\mathrm{e}\mathrm{V}$ relative to DFT. This demonstrates that the model maintains a level of accuracy comparable to that achieved for CsPbBr3.
@@
Figure Application to additional compositions with experimental benchmarksa compares the band gaps of MAPbBr3 obtained from DFT and Hamster across different supercell sizes at $300\text{\,}\mathrm{K}$. Leveraging the computational efficiency of Hamster, we extend our simulations to $16\times 16\times 16$ supercells comprising almost 50,000 atoms. The predicted band gap of MAPbBr3 decreases by approximately $0.2\text{\,}\mathrm{e}\mathrm{V}$ as the supercell size increases, in line with a recent report on FAPbI3 [44]. Agreement between Hamster and DFT is observed up to $4\times 4\times 4$ supercells (768 atoms), beyond which direct DFT calculations become impractical.
@@
Finally, Fig. Application to additional compositions with experimental benchmarksb presents the variation of the band gap with temperature as predicted by Hamster across different supercell sizes, together with experimental data for comparison. The objective here is not to reproduce the experimental trend with exact quantitative accuracy, but to employ it as a benchmark for assessing the reliability of the model. Although minor deviations between results obtained with Hamster and experimental data remain, the predicted temperature dependence captures the experimental behavior well, with the agreement becoming more favorable at larger supercell sizes. This consistency between experiment and computation supports the robustness and overall validity of the approach.
@@
Fig. 4 : Band gap of MAPbBr3 as a function of supercell size and temperature. a Band gap computed with Hamster (blue) and DFT (black) for varying supercell sizes at a temperature of $300\text{\,}\mathrm{K}$. The $x$-axis indicates the supercell dimension $n$, corresponding to an $n\times n\times n$ replication of the cubic cell. b Temperature-induced change in band gap, $\Delta E_{\mathrm{g}}$, references to the band-gap value at a temperature of $300\text{\,}\mathrm{K}$, computed for varying supercell (green: $4\times 4\times 4$; dark orange: $16\times 16\times$) sizes with Hamster and compared to an experimentally-determined fit function (grey) [45].

### v2

@@ Introduction
Relevant in this context, learning the Hamiltonian is challenging because it depends not only on a system’s atomic structure but also on the chosen representation, so there is no single unique mapping from structure to matrix elements. However, equivariant ML models respecting fundamental physical transformations, such as rotations and translations, ensure that predictions transform consistently under symmetry operations, and thereby significantly improve data efficiency and generalization. Specifically, equivariant neural networks (ENNs) were recently developed as particularly promising frameworks for Hamiltonian learning [17, 18, 19]. Although ENNs improve data efficiency by encoding symmetry and therefore require fewer samples than comparable non-equivariant models, they may still demand substantial amounts of training data to capture complex dynamical behavior, such as that exhibited in multi-component materials at finite temperature. Related to this limitation, existing ENNs for Hamiltonian learning usually rely on access to the real-space Hamiltonian represented in a consistent basis, which restricts their applicability to specific DFT frameworks requiring careful benchmarking [20].
@@ Results
Fig. 1 : Overview of the physics-informed Hamiltonian learning model and workflow with results for GaAs. a Illustration of pairwise and environment-dependent electronic interactions arising from structural fluctuations in a chemical system and their effect on the electronic structure captured via an effective Hamiltonian, $\hat{H}_{\mathrm{eff}}$, in the Hamster approach. It combines a physical model and environment-dependent machine learning terms. b Schematic visualization of the environment descriptor for the matrix element between atoms $i$ and $j$, which treats local environments of the two atoms separately. Different atomic species are indicated by purple and gray colors. Atoms within a distance $r_{\mathrm{cut}}$, which is indicated by dashed circles, are labeled as $k_{x}$, with $x=1,2,3,..$. The $s$ and $p$ orbitals of selected atoms are shown schematically as red circles and red-blue handles, respectively. c Validation error, calculated as mean absolute error (MAE) across all eigenvalues, for a Hamster model trained on an increasing number of training structures for GaAs at $400\text{\,}\mathrm{K}$. d Comparison of residuals (difference between model and DFT eigenvalues) of a pristine tight binding (orange) and the Hamster model (blue) with respect to density functional theory (DFT) data, averaged over k-points and 100 snapshots, at $400\text{\,}\mathrm{K}$. Dashed vertical lines indicate the valence band maximum (VBM).
@@
We employ a radial basis function kernel model, where each kernel support point, $x_{n}$, is associated with a learnable parameter, $h_{n}$. Given a trial environment descriptor vector, $\textbf{x}^{\textbf{R}}_{ij}$, the ML correction for the respective matrix element in Eq. 3 is calculated as
[…]
@@
where the sum runs over all atoms and orbitals, indexed by $k$, within a cutoff radius, $r_{\mathrm{cut}}$. The smooth cutoff function, $f_{\mathrm{cut}}(r)=\frac{1}{2}[\cos(\frac{\pi r}{r_{\mathrm{cut}}})+1]$, ensures a gradual decay of contributions from distant atoms. The local environment is identical for all lattice translation vectors, R, as a consequence of translational symmetry. Note that the integrals in Eq. 5 reduce to those in Eq. 3 when $V_{v}=1$, and therefore need to be evaluated only once for both the TB and ML contributions.
To provide the model with information about the interaction type, we augment the descriptor vector with several structural features: the equilibrium distance between atoms $i$ and $j$, the angular orientation of their respective orbitals relative to the connecting vector, $\textbf{r}_{ij}$, and the atomic numbers $Z_{i}$ and $Z_{j}$. These quantities can either be kept fixed during atomic motion or updated dynamically, thereby providing the model with the flexibility to accommodate additional physical effects such as thermal lattice expansion. This design ensures that the ML model focuses on learning changes in the local environment, thereby improving data efficiency by reducing the complexity of the learning task. The descriptor vector then reads
$x^{\textbf{R}}_{ij}=(Z_{i},Z_{j},\Delta r^{\textbf{R}}_{ij},\theta^{\textbf{R}}_{i},\theta^{\textbf{R}}_{j},\varphi^{\textbf{R}}_{ij},h_{\mathrm{env},i},h_{\mathrm{env},j})^{\mathrm{T}}.$ (6)
Here, $\Delta r^{\textbf{R}}_{ij}$ is the distance between atoms $i$ and $j$ with the translation vector R applied to atom $j$, $\theta^{\textbf{R}}_{x}$ ($x=i,j$) is the angle between the orientation of the orbital on atom $x$ and the bonding axis, and $\varphi^{\textbf{R}}_{ij}$ is the angle between the orbital orientations on atoms $i$ and $j$.
Finally, we emphasize that the ordering of values within the descriptor vector in Eq. 6 is critical, as inconsistent arrangements can violate the symmetry properties of the Hamiltonian. To preserve these symmetries, the descriptor is constructed such that its structure remains invariant under operations that should produce equal matrix elements, for example the relation $H^{\textbf{R}}_{ij}=H^{-\textbf{R}}_{ji}$ (see, e.g., Ref. 16 for a detailed discussion of Hamiltonian symmetry properties). This should be distinguished from equivariance, which instead ensures that the model output transforms consistently under symmetry operations, as employed in equivariant models [15, 31, 18, 13, 14].
@@
Fig. 2 : Training and descriptor analysis for CsPbBr3. a Dependence of training (dark orange) and validation (gray) loss (MAE between DFT and Hamster energy eigenvalues) on the number of training structures. The data point at zero training structures corresponds to the underlying TB model without ML corrections. b Principal component analysis of the kernel support points for 12 training structures. In both panels, we compare two different cutoff radii ($r_{\mathrm{cut}}$) for the interaction range, namely 5.5 $5\text{\,}\text{\AA }$ and $6.2\text{\,}\text{\AA }$.
Fig. 3 : Transferability of the model across temperatures and large-scale calculations. a Band gap of CsPbBr3 computed with DFT (black) and Hamster (blue) for a $2\times 2\times 2$ supercell of CsPbBr3. Results obtained using Hamster for a $16\times 16\times 16$ supercell are shown as well (green). Thin vertical lines indicate the standard deviations. The gray line represents the slope of the fitting function derived from experimental data [40]. b Comparison of residuals of a pristine TB (orange) and the Hamster model (blue) with respect to DFT data, averaged over $k$-points and 100 snapshots, for CsPbBr3 at $425\text{\,}\mathrm{K}$.
@@
The learning curves (see Fig. Physics-informed Hamiltonian learning for halide perovskitesa) show that the validation error plateaus after using around 10 training structures, reaching an accuracy for the eigenvalues that drops below $50\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$. While this level of accuracy is comparable to that reported in a recent study [21] using a neural-network scheme for elemental and binary semiconductor compounds, we emphasize that our approach needs far fewer training structures for the considerably more complex halide perovskite CsPbBr3, well below the few hundred typically required by neural-network-based Hamiltonian learning models [17, 31, 21]. Remarkably, for this material Hamster still generalizes reasonably well with as few as two training structures, remaining within $0.03\text{\,}\mathrm{e}\mathrm{V}$ of the training error, highlighting the markedly superior data efficiency of our approach. The mild overfitting at very low training set sizes, notable through the small gap between training and validation error, disappears as the training set increases.
@@
We analyze the structure of the descriptor space to gain insight into the model’s increased expressiveness when a larger $r_{\mathrm{cut}}$ is used. To this end, we perform a principal component analysis (PCA, see Methods) on the sampled descriptors for both cutoff radii, which projects the eight-dimensional descriptor vectors onto a two-dimensional subspace while preserving as much of their variance as possible. This enables a clear visualization of their distribution (see Fig. Physics-informed Hamiltonian learning for halide perovskitesb). In this representation, descriptor vectors that encode types of interactions lie close together. Since each descriptor component corresponds to a specific orbital interaction in the Hamiltonian (see Eq. 6), the PCA projection groups and separates regions of descriptor space according to the orbital couplings they represent.
@@
Before turning to large-scale simulations, we apply the model to MD trajectories of a $2\times 2\times 2$ supercell of CsPbBr3 for benchmarking against DFT data, using $r_{\mathrm{cut}}=$ 5.5 $5\text{\,}\text{\AA }$ in our model. Fig. Physics-informed Hamiltonian learning for halide perovskitesa compares the temperature-dependence of the band gap in CsPbBr3 as obtained with Hamster and DFT. Although the model was trained solely on structures from an MD trajectory at $425\text{\,}\mathrm{K}$, its predictions remain in close agreement with DFT also at higher temperatures. Across all three considered temperatures, the difference between the DFT-computed band gap and the value predicted by Hamster remains below $50\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$, and the MAEs of energy eigenvalues are $0.047\text{\,}\mathrm{e}\mathrm{V}$ ($425\text{\,}\mathrm{K}$), $0.049\text{\,}\mathrm{e}\mathrm{V}$ ($525\text{\,}\mathrm{K}$), and $0.057\text{\,}\mathrm{e}\mathrm{V}$ ($625\text{\,}\mathrm{K}$). Furthermore, we find that the fluctuations of instantaneous band gaps, gauged by the standard deviation at each temperature, are similar in DFT and Hamster. These results underscore the transferability of our model over a considerable temperature range within the cubic phase of CsPbBr3.
@@
A key advantage of our approach is that it enables simulations of system sizes far beyond the practical reach of explicit first-principles methods such as DFT. Modeling of larger systems can reveal distinct dynamical behavior and, by reducing finite-size effects, provide a more realistic description of the material. To explore this regime, we train an MLFF for CsPbBr3 (see Methods) and perform MD simulations for a $16\times 16\times 16$ supercell of CsPbBr3 consisting of 20480 atoms. Using the Hamster model, we subsequently predict Hamiltonians and compute band gaps by extracting eigenvalues at the VBM and CBM via the Lanczos algorithm as implemented in Arpack [44]. Figure Physics-informed Hamiltonian learning for halide perovskitesa shows that the mean value of the band gap at $425\text{\,}\mathrm{K}$ closely matches the one obtained for the smaller supercell. At the same time, the sizable band-gap fluctuations seen for the smaller supercells are almost entirely averaged out, yielding a much narrower distribution and a more realistic description of the material at finite temperature. This improved statistical convergence allows us to assess the temperature dependence of the band gap. While the absolute value of the band gap is affected by the known inaccuracies of PBE-level DFT, the resulting temperature dependence can still be meaningfully compared to experiment. For the $16\times 16\times 16$ supercell, we obtain a slope of $3.95\text{\times}{10}^{-4}\text{\,}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{K}$, in good agreement with the experimental value of $3.41\text{\times}{10}^{-4}\text{\,}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{K}$ reported in Ref. [40]. We note, however, that thermal expansion is not included in the simulation and that the temperature range considered here extends beyond that of the experimental study, both of which are expected to affect the band gap values.
@@
We optimize the model parameters following the same procedure as for CsPbBr3 (see Methods section), again using $r_{\mathrm{cut}}=$ 5.5 $5\text{\,}\text{\AA }.$ However, for this material we do not include orbitals stemming from the cation and perform classical force-field MD calculations (see Methods section). When applied to an independent test set consisting of 100 MD snapshots of a $2\times 2\times 2$ supercell at $300\text{\,}\mathrm{K}$, the model reaches an MAE of $0.060\text{\,}\mathrm{e}\mathrm{V}$ relative to DFT. This demonstrates that the model maintains a level of accuracy comparable to that achieved for CsPbBr3.
@@
Figure Application to additional compositions with experimental benchmarksa compares the band gaps of MAPbBr3 obtained from DFT and Hamster across different supercell sizes at $300\text{\,}\mathrm{K}$. Leveraging the computational efficiency of Hamster, we extend our simulations to $16\times 16\times 16$ supercells comprising almost 50,000 atoms. The predicted band gap of MAPbBr3 decreases by approximately $0.15\text{\,}\mathrm{e}\mathrm{V}$ as the supercell size increases, in line with a recent report on FAPbI3 [45]. Agreement between Hamster and DFT is observed up to $4\times 4\times 4$ supercells (768 atoms), beyond which direct DFT calculations become impractical.
@@
Figure Application to additional compositions with experimental benchmarksb presents the variation of the band gap with temperature as predicted by Hamster across different supercell sizes, together with experimental data for comparison. The objective here is not to reproduce the experimental trend with exact quantitative accuracy, but to employ it as a benchmark for assessing the reliability of the model. Although minor deviations between results obtained with Hamster and experimental data remain, the predicted temperature dependence captures the experimental behavior well, with the agreement becoming more favorable at larger supercell sizes. This consistency between experiment and computation supports the robustness and overall validity of the approach.
@@
Fig. 4 : Band gap of MAPbBr3 as a function of supercell size and temperature, with associated computational scaling. a Band gap computed with Hamster (blue) and DFT (black) for varying supercell sizes at a temperature of $300\text{\,}\mathrm{K}$. The $x$-axis indicates the supercell dimension $n$, corresponding to an $n\times n\times n$ replication of the cubic cell. b Temperature-induced change in band gap, $\Delta E_{\mathrm{g}}$, references to the band-gap value at a temperature of $300\text{\,}\mathrm{K}$, computed for varying supercell sizes (green: $4\times 4\times 4$; dark orange: $16\times 16\times 16$) with Hamster and compared to an experimentally-determined fit function (gray) [40]. c Runtime of Hamster evaluated on 10 structures as a function of cell size. Timings are shown for full diagonalization of the Hamiltonian (full diag; green stars), partial diagonalization of six eigenvalues around the VBM using the Lanczos algorithm (partial diag; orange diamonds), and construction and saving of the Hamiltonian without diagonalization (no diag; blue circles). A constant Julia compilation time of 44 $4\text{\,}\mathrm{s}$ has been subtracted. Reference lines indicating linear and cubic scaling are shown in gray.
To assess the computational performance of Hamster, we benchmark the total runtime for three calculation modes: (i) full diagonalization of the Hamiltonian, (ii) Lanczos-based partial diagonalization targeting six eigenvalues near the VBM, and (iii) construction and output of the Hamiltonian matrix in Eq. 1 without solving the eigenvalue problem. For each benchmark, the model is evaluated for 10 independent structures, using the same configurations as in Fig. Application to additional compositions with experimental benchmarks, and all calculations are performed on a single compute node using two Intel Xeon Platinum 8168 CPUs. Full diagonalization of the largest system size ($16\times 16\times 16$) is not tractable on the chosen, rather modest hardware and is therefore omitted. The resulting runtimes are shown in Fig. Application to additional compositions with experimental benchmarksc. Since the benchmarks include the first invocation of each function, we fit a function, $T(x)=T_{0}+ax^{n}$, to the timings and subtract the constant time $T_{0}=$44\text{\,}\mathrm{s}$\$, i.e., the initial Julia compilation overhead, which disproportionately affects benchmarks with short execution times.
The observed scaling behavior is consistent with theoretical expectations. The construction of the Hamiltonian exhibits nearly ideal linear scaling with system size. Partial diagonalization remains significantly faster than full diagonalization, but deviates from linear behavior, showing an effective scaling exponent of approximately $1.8$. This deviation originates from the shift–invert transformation, which can reduce the sparsity of the Hamiltonian. In contrast, full diagonalization yields a scaling exponent of $2.8$, close to the cubic scaling expected for conventional DFT implementations. It is worth noting that the absolute computational cost of DFT calculations is substantially higher, primarily due to the overhead associated with repeated self-consistent field iterations and the much larger underlying basis sets. Additional speedups for diagonalization could be achieved through more advanced algorithms, highly parallel implementations, or GPU-accelerated solvers.
Within our approach, the wall time for the largest system considered remains below one hour, demonstrating that Hamster can efficiently handle large-scale systems even on modest hardware. When only the Hamiltonian is required and eigenvalues are not computed, genuinely linear scaling can be achieved. Moreover, many important physical quantities, such as transport properties or the density of states, can be evaluated with linear-scaling methods that exploit the sparsity of the Hamiltonian [46, 47]. Combining our Hamiltonian learning framework with such techniques therefore provides a promising route toward efficient simulations of large-scale systems.